OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDALS!!! For the system of linear equations below, use graphing to determine in which quadrant the solution lies. 3x+4y=-12 -7x+5y=-35 Quadrant I Quadrant II Quadrant III Quadrant IV

OpenStudy (pottersheep):

So graph it! Do you know how? Hint: Put it into y=mx+b form

OpenStudy (pottersheep):

3x+4y=-12 4y = -3x-12 y=-3/4x - 3

OpenStudy (pottersheep):

-7x+5y=-35 5y=7x-35 y=7/5x - 7

OpenStudy (anonymous):

Im terrible at graphing, what is the answer?

OpenStudy (anonymous):

@anonymous_user @Abmon98 @mathmath333 @mathgeek27

OpenStudy (anonymous):

@pottersheep